Amber purchased Hampton Industries Inc. stock for $18.35 and sold it 6 months later for $21.45 after receiving a $0.50 dividend. What was her holding period return (HPR), Annual Percentage Rate (APR) and Effective Annual Rate (EAR)?
a) 19.62% ; 39.24% ; 43.09%
b) 20.34% ; 40.68% ; 9.7%
c) 14.17% ; 28.34% ; 30.35%
d) 20.34% ; 40.68% ; 44.82%
Answer #1

Holding period return = (Dividend + selling price - buying price)/buying price

= (0.5+21.45-18.35)/18.35

= 19.62%

APR = 19.62%*12/6 = 39.24%

Effective Annual Rate = (!+periodic Rate)^2 - 1

= (1+0.1962)^2 - 1

= 43.09%

Hence, the answer is a.